OpenStudy (anonymous):

. A 22 kg sled is pushed for 5.2 m with a horizontal force of 20 N, starting from rest. Ignore friction. Find the final speed of the sled.

OpenStudy (anonymous):

2.2 m/s 3.1 m/s 4.7 m/s 9.5 m/s

OpenStudy (anonymous):

work done = force x distance, so 5.2m x 20 = 104J, so thats kinetic energy, knowing mass we can use formula to get velocity so E = 1/2mv^2 so 104 = 1/2 * 22kg * v^2 rearrange for v = SQRT(208/22)

OpenStudy (anonymous):

which is approx 3.1m/s